Phindile buys two shirts at a clothing shop and each shirt cost R250 . She also buys three pairs of jeans for R450 each , and receives a 25% (112.50) discount on jeans bought. What is the amount did phindile pay
1
Expert's answer
2020-06-15T17:54:40-0400

cost of one shirt = 250

number of shirts bought = 2

total cost of shirts = 2"\times" 250 = 500

price of one jeans = 450

discount on one jeans = 112.50

actual cost of one jeans = 450-112.50=337.5

number of jeans bought = 3 pairs = 6

total cost of jeans = 6 "\times" 337.5 = 2025

Amount paid = 2525